OpenStudy (anonymous):

can you help? Use guess, check and comment to find two numbers whose sum and product are given. Sum = 83, product = 1632

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics factorize 1632 you get 1,2,2,2,2,2,51 we have to make a sum of 83 using these. which happens to be 32 and51 so the numbers are 32 and51

OpenStudy (anonymous):

x+y=82 xy=1632 x(82-x)=1632

OpenStudy (anonymous):

mrjoselson's first two equations when solved for x and y give: \[\{x\to 34\},\{x\to 48\} \] 34*48 = 1632